OpenStudy (anonymous):

If a species contains 23% A in its DNA, what is the percentage of guanine it would contain?

OpenStudy (blues):

Edwin Chargaff discovered that A and T are present in equal quantities, and C and G are present in equal quantities. Is that much clear?

OpenStudy (anonymous):

A=T and G=C or 2G then 100-A+T = 2G 100-23+23=2G 100-46=2G 54=2G G=27%